glen can paint a house is 9 hours while ced can paint the same house in 16 hours. they work together for 4 hours.after 4hours, ced left and len finished the job alone. how many hours did glen take to finish the job?
Answer by ptaylor(2198)


Let x=number of hours it took Glen to finish the job alone
Glen paints at the rate of 1/9 house per hour
Ced paints at the rate of 1/16 house per hour
Together they paint at the rate of 1/9 + 1/16=(16+9)/144 = 25/144 house per hour
In 4 hours, working together, they paint (25/144)*4=25/36 of the house leaving
1-25/36=11/36 of the house yet to be painted
So our equation to solve is:
1/9 of the house per hour times x hours equals 11/36 of the house, or
(1/9)*x=11/36 multiply each side by 36
4x=11
x=2 3/4 hours----time it takes Glen to finish
Hope this helps----ptaylor